Extra Credit Problem o Suppose a firm is expected to increase dividends by 20% for the first three years. After tha, dividends will increase at a rate of % per year indefinitely. If the most recent dividend was $2 and the required return is 1590, what isthe price of the stock?
Answer #1

Next year dividend = Current year dividend (1 + growth rate)

D1 = 2 (1 + 0.20) = 2 (1.2) = $2.40

D2 = 2.4 (1 + 0.20) = 2.4 (1.2) = $2.88

D3 = 2.88 (1 + 0.20) = 2.88 (1.2) = $3.456

D4 = 3.456 (1 + .05) = 3.456 (1.05) = $3.6288

Expected future price:

P3 = D4 / (r - g) = 3.6288 / (0.15 - 0.05) = $36.288

Present value of future cash flow:

P0 = (2.4 / 1.151) + [(2.88 + 36.288) / (1.152)] + [(3.456 + 36.288) / (1.153)]

= 2.0869565217 + (39.168 / 1.152) + (39.744 / 1.153)

= 2.0869565217 + 29.6166351607 + 26.1323251418

= $57.84 Ans.